# Will LLMs Scaling Hit the Wall? Breaking Barriers via Distributed Resources on Massive Edge Devices

### v1

@@ 3.2 Massive Computing Power from Edges
In recent years, edge computing has experienced explosive growth in computing power, driving smart devices to evolve from single-function tools into multimodal perception and decision-making centers. For instance, as shown in Figure 5, flagship smartphones such as the iPhone 16 series, equipped with 3nm process chips, have achieved computing power exceeding 2 TFLOPS (Apple Inc., 2024), enabling local execution of complex AI tasks like real-time image enhancement and multilingual speech translation. Notably, the computing power of an individual smartphone has potentially surpassed that of laptops in the same period. The breakthroughs are even more pronounced in the desktop sector, achieving an annual computing power growth rate of 1.29$\times$/year, surpassing that of smartphones and laptops (both 1.20$\times$/year). These three types of devices form a differentiated growth hierarchy, collectively driving edge computing’s overall computing power to expand at an annual average rate of 1.28$\times$/year. This growth is fueled by three key technological drivers: advanced manufacturing processes (3nm technology increases transistor density by 60% (Apple Inc., 2024)), dedicated architectures (modern smartphone SoCs integrate NPUs for AI acceleration), and scenario-driven innovation (e.g., autonomous driving demands end-to-end latency of less than 100ms (NVIDIA, 2023)).
@@
Edge computing has potential for LLM training.
We analyzed the performance of smartphone chips, representing typical edge devices, and estimated their overall computing power. To ensure our estimation is as accurate as possible, we based our calculations on the market share data from Canalys (2025). We then estimated the total computing power of newly produced mobile devices by averaging the chip performance of each brand. From 2020 to 2024, smartphone chip performance has seen significant improvements, with peak computing power increasing from 1.53 TFLOPS to 4.95 TFLOPS, and average computing power rising from 0.48 TFLOPS to 1.38 TFLOPS. Meanwhile, the overall computing power of mobile devices has grown from 817 EFLOPS in 2020 to 2,758 EFLOPS in 2024, and totally 9278 EFLOPS for past 5 years. This trend highlights the rapid expansion of edge computing power, which is not only essential for AI applications but also holds the potential for training complex AI models. For instance, training the DeepSeek-v3 (Liu et al., 2024a) model utilizes 2048 H100 GPUs, each providing a peak FP32 performance of 59.30 TFLOPS, resulting in a total computational capacity of 121,446.4 TFLOPS. If this workload were distributed across edge devices with a peak performance of 2 TFLOPS (e.g., mobile chips like the iPhone 16 series), approximately 60,723 users with edge devices working (ideally) in parallel would be required to match the computational capacity.
@@ 4.1 Small Language Models at Edges
The first move of AI to Edge is to deploy small language models (SLMs) to edge devices (Lu et al., 2024; Wang et al., 2024a; Van Nguyen et al., 2024). This trend is driven by the growing demand for AI applications that can run directly on edge devices, motivated by needs for privacy, offline usage, and real-time processing without cloud dependence. However, edge devices have limited memory, computation, and energy resources, requiring more efficient and compact models.
@@
The classic Transformer architecture (Vaswani et al., 2017) uses self-attention mechanisms for effective sequence modeling but faces quadratic complexity challenges, with models like TinyBERT (Jiao et al., 2020) (14.5M parameters) and ALBERT (Lan et al., 2020) (12M parameters) demonstrating its effectiveness at small scales. Mamba (Gu & Dao, 2023), based on state space models, achieves linear complexity and faster inference by utilizing only the previous hidden state, as demonstrated by Zamba2-2.7B (Glorioso et al., 2024) which achieves twice the speed and 27% reduced memory overhead compared to traditional models. Hymba (Dong et al., 2024) combines both approaches by integrating attention and SSM heads within the same layer for parallel processing, with its 1.5B variant trained on DCLM-Baseline-1.0 and SmolLM-Corpus achieving 11.67 times cache size reduction while outperforming Llama-3.2-3B. The xLSTM architecture (Beck et al., 2024) modernizes LSTM with exponential gates and matrix memory cells, with models ranging from 125M to 1.3B parameters trained on 300 billion tokens from SlimPajama (Systems & Computer, 2023), consistently outperforming comparable RWKV-4 (Peng et al., 2023), Llama (Inan et al., 2023), and Mamba models across various tasks in the PALOMA benchmark (Magnusson et al., 2023). These architectural innovations demonstrate the potential for efficient and powerful language models that can run effectively on edge devices.
@@
The construction of efficient SLMs relies on a comprehensive suite of techniques, each with specific performance trade-offs. For training SLMs from scratch, optimized MLM approaches (Devlin et al., 2019) with increased masking ratio (25% vs traditional 15%) demonstrate 2-3% performance improvements for models under 3B parameters. When deriving SLMs from existing LLMs, knowledge distillation has proven particularly effective, with response-based distillation (Sanh et al., 2019) reducing model size by 40% while maintaining 95% of the original performance. Similarly, model compression techniques like 4-bit quantization (Zhang et al., 2022) can reduce model size by 75% with only 2-3% performance drop. In architecture optimization, the Mixture of Experts approach (Fedus et al., 2022) enables 65% parameter reduction while potentially improving performance in specific tasks. Domain specialization has shown remarkable results, particularly in the medical field where 3B parameter models achieve 92% accuracy, outperforming 175B models (89%) (Fu et al., 2023). The combination of these techniques yields impressive results - a notable example is a 770M parameter model from (Shridhar et al., 2023) that combines distillation, quantization, and domain specialization to achieve 95% of the performance of a 540B model on specific tasks while requiring less than 0.15% of the computational resources. Most successful SLMs achieve their efficiency by combining multiple techniques, typically starting with knowledge distillation, applying compression methods, and finishing with domain-specific fine-tuning. Wang et al. (2024a) has provided a comprehensive survey of SLMs, and we summarize the architecture innovations and training methods in Appendix E.
@@ 4.2 Collaborative Inference at Edges
The emergence of collaborative inference at the edge represents a significant shift in AI infrastructure, driven by the increasing need for accessible and cost-effective AI inference solutions. Traditional large-scale data centers, while powerful, often present barriers in terms of cost, energy consumption, and accessibility, providing a middle ground between edge devices and massive data centers.
Exo (Labs, 2025) is a recently popular collaborative inference framework that enables users to create AI clusters using everyday devices such as iPhones, iPads, Android devices, Macs, and Linux systems. By leveraging a peer-to-peer architecture, Exo allows these devices to work together seamlessly, effectively unifying their computational resources into a powerful AI cluster. Neurosurgeon (Kang et al., 2017) introduces a lightweight scheduler that partitions deep neural network (DNN) computations between devices and datacenters, achieving significant reductions in latency and energy consumption. MoE2 (Jin et al., 2025) proposes a Mixture-of-Edge-Experts framework to optimize inference for large language models under energy and latency constraints. Edgent (Li et al., 2018a) enables low-latency edge intelligence through adaptive DNN partitioning and early-exit mechanisms. Galaxy (Ye et al., 2024) leverages a hybrid model parallelism approach and heterogeneity-aware planning to efficiently execute Transformer models across edge devices, reducing inference latency. These approaches collectively emphasize the synergy of device-edge collaboration for efficient deep learning inference in edge AI systems.
While collaborative inference at edge devices has made some strides, the current landscape has yet to fully capitalize on the immense potential of distributed data resources at the edge. The next frontier of research lies in transforming these devices from mere inference endpoints into active participants in the training process. This paradigm shift promises to usher in a new era of distributed AI that effectively harnesses the collective computing power and diverse data resources of billions of edge devices worldwide, fundamentally reshaping how we approach AI system development and deployment.
@@ 4.3 Collaborative Training at Edges
Federated learning (McMahan et al., 2017b) is a practical paradigm that enables collaborative model training while preserving data privacy. Instead of collecting raw data from edge devices, which may violate privacy regulations like GDPR (Parliament & of the European Union, 2016), this approach distributes the training process across multiple devices. Each device trains on its local data and only shares model updates with the central server. This approach can effectively utilizes both computational and data resources available at edges.
@@
Federated LLMs for fine-tuning has emerged as a critical direction in recent research of large language models, addressing the challenges of privacy preservation and resource constraints. FedLLM (Wu et al., 2024c) introduces a novel framework that enables efficient fine-tuning of large language models in a federated setting, demonstrating comparable performance to centralized fine-tuning while maintaining data privacy. FedFM (Chen et al., 2024b) tackles the critical challenges of system and statistical heterogeneity in federated learning, proposing adaptive optimization techniques that improve model convergence across diverse client devices. To address the computational constraints of edge devices, FedPET (Li et al., 2024e) employs parameter-efficient fine-tuning techniques, significantly reducing the memory and computation requirements while maintaining model performance. This approach enables even resource-constrained devices to participate in the fine-tuning process. Similarly, (Xu et al., 2024) bridges the gap between federated learning and foundation models by introducing novel techniques for efficient knowledge transfer and model adaptation in distributed settings. In domain-specific applications, FedQA (Wang et al., 2024b) demonstrates the effectiveness of federated learning for question-answering tasks, showing that models can be fine-tuned on sensitive domain-specific data while preserving privacy. These advancements are supported by open-source frameworks like Flower (Flowerlab, 2025) and FATE-LLM (Fan et al., 2023b), which provide robust platforms for implementing federated fine-tuning of large language models.
@@
Federated LLMs for pretraining have opened up exciting new possibilities for large language model training. Rather than relying on traditional data center approaches, researchers have developed innovative geographically distributed frameworks that enable collaborative training across many devices. Notably, Prime Intellect (PrimeIntellect-ai, 2025) has launched the first decentralized training project for a 10 billion parameter model, named INTELLECT-1, which utilizes the OpenDiLoCo framework to significantly reduce communication costs between nodes. This innovative approach allows for dynamic management of computational resources across multiple locations, achieving an impressive 83% overall computational utilization while collaborating with up to 112 H100 GPUs across five countries and three continents. The model not only enhances parameter efficiency by 25 times but also demonstrates robust performance in various benchmark tests. In parallel, the Flower Lab (Flowerlab, 2025) has introduced FlowerLLM, which successfully trained a 1.3 billion parameter large language model (LLM) using novel federated learning methods. Additionally, it has also developed Photon (Sani et al., 2024), an open-source framework that provides flexible configurations for training models of different sizes, making federated LLM training more accessible to a broader range of participants and computational resources.
@@
5 An Open Problem: How to Train Large Language Models with Small Edge Devices?
A fundamental limitation of traditional federated learning lies in its requirement for each participant to maintain and train a complete model locally. This assumption becomes particularly problematic in the context of large language models, where the computational and memory requirements far exceed the capabilities of most individual participants. For instance, in sensitive domains like healthcare (Bolton et al., 2024), multiple hospitals may wish to collaboratively train a medical language model to leverage their collective data while maintaining data privacy. However, traditional federated learning mandates that each participating hospital possess sufficient computational resources to train the complete model locally. Despite these institutions’ valuable data contributions and their motivation to enhance model capabilities through collaborative training, many hospitals lack the necessary infrastructure to participate effectively. This resource constraint significantly limits the potential for collaborative model training in critical domains where data privacy is paramount but computational resources are unevenly distributed (Kairouz et al., 2021).
@@ 5 An Open Problem: How to Train Large Language Models with Small Edge Devices?
While some federated learning approaches allow training partial model parameters (Horvath et al., 2021; Alam et al., 2022), the enormous disparity between large language models and what edge devices can train—often orders of magnitude smaller due to inherently constrained resources—remains too vast to be effectively bridged by current FL frameworks. Therefore, we need to develop a new federated learning paradigm that enables participants to collaboratively train a large language model even under extremely limited resources (as shown in Figure 6). It is still an open problem to train large language models with small edge devices. Therefore, we encourage the research community to develop novel distributed collaborative computing methods in two key directions:
5.1 Heterogeneous Device Model Fusion:
from Small to Large
@@ 5.1 Heterogeneous Device Model Fusion:
The first direction addresses the fundamental challenge of model size disparity in federated learning. Modern large language models typically contain hundreds of billions of parameters, while edge devices have severely limited computational resources. This creates an enormous scale gap - the large target model may be hundreds or even thousands of times larger than what individual devices can handle. To bridge this gap, each edge device should run a small language model that matches its computational capacity. For example, while the central model may have 100 billion parameters, a resource-constrained mobile device might only handle a 100-million parameter model, representing a 1000x size difference. The key challenge then becomes how to effectively aggregate and fuse knowledge from these much smaller models into the large target model. We need novel techniques that can meaningfully combine insights from models operating at radically different scales while preserving the unique contributions of each small model. This requires fundamentally rethinking traditional model fusion approaches (Velasevic et al., 2023; Azizan-Ruhi et al., 2019) to handle such extreme parameter count disparities.
5.2 Heterogeneous Device Compute Sharing:
from Node to Cluster
@@ 5.2 Heterogeneous Device Compute Sharing:
The second direction is to enable efficient compute resource sharing across heterogeneous devices by treating them as a unified compute cluster rather than independent nodes. Consider a smart home environment where multiple devices—smartphones, laptops, and desktop computers—could form a collaborative compute cluster. While each individual device has limited resources, their collective computing power could be substantial. For example, a laptop could handle intensive computational tasks, smartphones could manage coordination and lightweight processing, and desktop computers could contribute their onboard computing power. Meanwhile, other IoT devices such as smart speakers, security cameras, and vehicles could serve as data sources, providing valuable real-world inputs like voice commands, visual feeds, and environmental parameters. The language model would effectively run and train across this entire device cluster, leveraging both computing power and diverse training data from the environment. This distributed execution requires new frameworks that can intelligently decompose and distribute model computations based on each device’s capabilities and current load. The system must dynamically balance workloads - when the security cameras are idle at night, they could take on additional compute tasks, while during peak usage hours, the load could shift to other devices. This requires innovations in real-time resource allocation, task scheduling across heterogeneous hardware, and efficient inter-device communication protocols to ensure the collective computing power is optimally utilized (Zhao et al., 2024).

### v2

@@ 3.2 Massive Computing Power from Edges
In recent years, edge computing has experienced explosive growth in computing power, driving smart devices to evolve from single-function tools into multimodal perception and decision-making centers. For instance, as shown in Figure 5, flagship smartphones such as the iPhone 16 series, equipped with 3nm process chips, have achieved computing power exceeding 2 TFLOPS [52], enabling local execution of complex AI tasks like real-time image enhancement and multilingual speech translation. Notably, the computing power of an individual smartphone has potentially surpassed that of laptops in the same period. The breakthroughs are even more pronounced in the desktop sector, achieving an annual computing power growth rate of 1.29$\times$/year, surpassing that of smartphones and laptops (both 1.20$\times$/year). These three types of devices form a differentiated growth hierarchy, collectively driving edge computing’s overall computing power to expand at an annual average rate of 1.28$\times$/year. This growth is fueled by three key technological drivers: advanced manufacturing processes (3nm technology increases transistor density by 60% [52]), dedicated architectures (modern smartphone SoCs integrate NPUs for AI acceleration), and scenario-driven innovation (e.g., autonomous driving demands end-to-end latency of less than 100ms [53]).
Figure 4: Edge Computing Power Evolution Trend. (Data source: [54]).
Figure 5: Smartphone Market Share and Computing Power Trends. (Data source: [55]).
Edge computing has potential for LLM training.
We analyze the performance of smartphone chips, representing typical edge devices, and estimated their overall computing power. To ensure our estimation is as accurate as possible, we based our calculations on the market share data from [55]. We then estimated the total computing power of newly produced mobile devices by averaging the chip performance of each brand. From 2020 to 2024, smartphone chip performance has seen significant improvements, with peak computing power increasing from 1.53 TFLOPS to 4.95 TFLOPS, and average computing power rising from 0.48 TFLOPS to 1.38 TFLOPS. Meanwhile, the overall computing power of mobile devices has grown from 817 EFLOPS in 2020 to 2,758 EFLOPS in 2024, and totally 9278 EFLOPS for past 5 years. This trend highlights the rapid expansion of edge computing power, which is not only essential for AI applications but also holds the potential for training complex AI models. For instance, training the DeepSeek-v3 [11] model utilizes 2048 H100 GPUs, each providing a peak FP32 performance of 59.30 TFLOPS, resulting in a total computational capacity of 121,446.4 TFLOPS. If this workload were distributed across edge devices with a peak performance of 2 TFLOPS (e.g., mobile chips like the iPhone 16 series), approximately 60,723 users with edge devices working (ideally) in parallel would be required to match the computational capacity.
@@ 4.1 Small Language Models at Edges
The first move of AI to Edge is to deploy small language models (SLMs) to edge devices [56, 57, 58]. This trend is driven by the growing demand for AI applications that can run directly on edge devices, motivated by needs for privacy, offline usage, and real-time processing without cloud dependence. However, edge devices have limited memory, computation, and energy resources, requiring more efficient and compact models.
@@
The classic Transformer architecture [39] uses self-attention mechanisms for effective sequence modeling but faces quadratic complexity challenges, with models like TinyBERT [59] (14.5M parameters) and ALBERT [60] (12M parameters) demonstrating its effectiveness at small scales. Mamba [61], based on state space models, achieves linear complexity and faster inference by utilizing only the previous hidden state, as demonstrated by Zamba2-2.7B [62] which achieves twice the speed and 27% reduced memory overhead compared to traditional models. Hymba [63] combines both approaches by integrating attention and SSM heads within the same layer for parallel processing, with its 1.5B variant trained on DCLM-Baseline-1.0 and SmolLM-Corpus achieving 11.67 times cache size reduction while outperforming Llama-3.2-3B. The xLSTM architecture [64] modernizes LSTM with exponential gates and matrix memory cells, with models ranging from 125M to 1.3B parameters trained on 300 billion tokens from SlimPajama [65], consistently outperforming comparable RWKV-4 [66], Llama [10], and Mamba models across various tasks in the PALOMA benchmark [67]. These architectural innovations demonstrate the potential for efficient and powerful language models that can run effectively on edge devices.
@@
The construction of efficient SLMs relies on a comprehensive suite of techniques, each with specific performance trade-offs. For training SLMs from scratch, optimized MLM approaches [68] with increased masking ratio (25% vs traditional 15%) demonstrate 2-3% performance improvements for models under 3B parameters. When deriving SLMs from existing LLMs, knowledge distillation has proven particularly effective, with response-based distillation [69] reducing model size by 40% while maintaining 95% of the original performance. In architecture optimization, the Mixture of Experts approach [70] enables 65% parameter reduction while potentially improving performance in specific tasks. Domain specialization has shown remarkable results, particularly in the medical field where 3B parameter models achieve 92% accuracy, outperforming 175B models (89%) [71]. The combination of these techniques yields impressive results - a notable example is a 770M parameter model from [72] that combines distillation, quantization, and domain specialization to achieve 95% of the performance of a 540B model on specific tasks while requiring less than 0.15% of the computational resources. Most successful SLMs achieve their efficiency by combining multiple techniques, typically starting with knowledge distillation, applying compression methods, and finishing with domain-specific fine-tuning. [57] has provided a comprehensive survey of SLMs, and we summarize the architecture innovations and training methods in Appendix E.
@@ 4.2 Collaborative Inference at Edges
The emergence of collaborative inference at the edge represents a significant shift in AI infrastructure, enabling more accessible and cost-effective AI solutions compared to traditional data centers which often present barriers in terms of cost, energy consumption, and accessibility. Recent frameworks like Exo [73] enable users to create AI clusters using everyday devices such as phones, tablets, and computers through peer-to-peer architecture, effectively unifying their computational resources.
Several approaches advance this paradigm: Neurosurgeon [74] introduces a lightweight scheduler that partitions DNN computations between devices and datacenters; MoE2 [75] optimizes LLM inference under energy and latency constraints; Edgent [76] enables low-latency edge intelligence through adaptive partitioning and early-exit mechanisms; and Galaxy [77] leverages hybrid model parallelism to efficiently execute Transformer models across edge devices. Despite these advances, the current landscape has yet to fully capitalize on the potential of distributed data resources at the edge. The next frontier lies in transforming these devices from mere inference endpoints into active training participants, representing a significant opportunity for distributed AI development.
4.3 Feasibility of On-Device Training
Recent advancements in optimization techniques have significantly reduced the memory and computational requirements for training machine learning models, making on-device training increasingly feasible, even on resource-constrained edge devices. For instance, Lin et al. [78] demonstrated training neural networks on microcontroller units with only 256KB of RAM by employing an algorithm-system co-design framework. Similarly, Cai et al. [79] introduced a tiny transfer learning approach that freezes most parameters and only trains a small subset, allowing effective fine-tuning with minimal memory requirements. Qiu et al. [80] proposed ZeroFL, a framework that relies on highly sparse operations to accelerate on-device training in federated learning settings, enabling efficient model training on edge devices with up to 95% sparsity. Recent work by Sugiura and Matsutani [81] further advanced this field with ElasticZO, which combines zeroth-order and first-order optimization to achieve a better trade-off between accuracy and training cost. Their ElasticZO-INT8 variant achieves integer arithmetic-only training, further reducing memory usage and training time by approximately 1.5x without compromising accuracy. These advancements suggest that on-device training is no longer limited to high-end devices with abundant resources. Even small embedded systems with memory capacities measured in kilobytes rather than gigabytes can participate in model training.
@@ 4.4 Collaborative Training at Edges
Federated learning [82] is a practical paradigm that enables collaborative model training while preserving data privacy. Instead of collecting raw data from edge devices, which may violate privacy regulations like GDPR [22], this approach distributes the training process across multiple devices. Each device trains on its local data and only shares model updates with the central server. This approach can effectively utilizes both computational and data resources available at edges.
@@
Federated LLMs for fine-tuning has emerged as a critical direction in recent research of large language models, addressing the challenges of privacy preservation and resource constraints. FlexLoRA [83] introduces a novel framework that enables efficient fine-tuning of large language models in a federated setting, demonstrating comparable performance to centralized fine-tuning while maintaining data privacy. FedFM [84] tackles the critical challenges of system and statistical heterogeneity in federated learning, proposing adaptive optimization techniques that improve model convergence across diverse client devices. To address the computational constraints of edge devices, FedPETuning [85] employs parameter-efficient fine-tuning techniques, significantly reducing the memory and computation requirements while maintaining model performance. This approach enables even resource-constrained devices to participate in the fine-tuning process. Similarly, [86] bridges the gap between federated learning and foundation models by introducing novel techniques for efficient knowledge transfer and model adaptation in multi-modal heterogeneous federated learning settings. In domain-specific applications, FedMatch [87] demonstrates the effectiveness of federated learning for question-answering tasks, showing that models can be fine-tuned on sensitive domain-specific data while preserving privacy. These advancements are supported by open-source frameworks like Flower [88] and FATE-LLM [89], which provide robust platforms for implementing federated fine-tuning of large language models.
@@
Federated LLMs for pretraining have opened up exciting new possibilities for large language model training. Rather than relying on traditional data center approaches, researchers have developed innovative geographically distributed frameworks that enable collaborative training across many devices. Notably, Prime Intellect [90] has launched the first decentralized training project for a 10 billion parameter model, named INTELLECT-1, which utilizes the OpenDiLoCo framework to significantly reduce communication costs between nodes. This innovative approach allows for dynamic management of computational resources across multiple locations, achieving an impressive 83% overall computational utilization while collaborating with up to 112 H100 GPUs across five countries and three continents. The model not only enhances parameter efficiency by 25 times but also demonstrates robust performance in various benchmark tests. In parallel, the Flower Lab [88] has introduced FlowerLLM, which successfully trained a 1.3 billion parameter large language model (LLM) using novel federated learning methods. Additionally, it has also developed Photon [91], an open-source framework that provides flexible configurations for training models of different sizes, making federated LLM training more accessible to a broader range of participants and computational resources.
@@
5 An Open Problem: How to Train LLMs with Small Edge Devices?
A fundamental limitation of traditional federated learning lies in its requirement for each participant to maintain and train a complete model locally. This assumption becomes particularly problematic in the context of large language models, where the computational and memory requirements far exceed the capabilities of most individual participants. For instance, in sensitive domains like healthcare [92], multiple hospitals may wish to collaboratively train a medical language model to leverage their collective data while maintaining data privacy. However, traditional federated learning mandates that each participating hospital possess sufficient computational resources to train the complete model locally. Despite these institutions’ valuable data contributions and their motivation to enhance model capabilities through collaborative training, many hospitals lack the necessary infrastructure to participate effectively. This resource constraint significantly limits the potential for collaborative model training in critical domains where data privacy is paramount but computational resources are unevenly distributed [93].
@@ 5 An Open Problem: How to Train LLMs with Small Edge Devices?
While some federated learning approaches allow training partial model parameters [94, 95], the enormous disparity between large language models and what edge devices can train—often orders of magnitude smaller due to inherently constrained resources—remains too vast to be effectively bridged by current FL frameworks. Therefore, we need to develop a new federated learning paradigm that enables participants to collaboratively train a large language model even under extremely limited resources (as shown in Figure 6). It is still an open problem to train large language models with small edge devices. Therefore, we encourage the research community to develop novel distributed collaborative computing methods in two key directions:
5.1 Heterogeneous Device Model Fusion: from Small to Large
@@
The first direction addresses the fundamental challenge of model size disparity in federated learning. Modern large language models typically contain hundreds of billions of parameters, while edge devices have severely limited computational resources. This creates an enormous scale gap - the large target model may be hundreds or even thousands of times larger than what individual devices can handle. To bridge this gap, each edge device should run a small language model that matches its computational capacity. For example, while the central model may have 100 billion parameters, a resource-constrained mobile device might only handle a 100-million parameter model, representing a 1000x size difference. The key challenge then becomes how to effectively aggregate and fuse knowledge from these much smaller models into the large target model. We need novel techniques that can meaningfully combine insights from models operating at radically different scales while preserving the unique contributions of each small model. This requires fundamentally rethinking traditional model fusion approaches [96, 97] to handle such extreme parameter count disparities.
5.2 Heterogeneous Device Compute Sharing: from Node to Cluster
@@
The second direction is to enable efficient compute resource sharing across heterogeneous devices by treating them as a unified compute cluster rather than independent nodes. Consider a smart home environment where multiple devices—smartphones, laptops, and desktop computers—could form a collaborative compute cluster. While each individual device has limited resources, their collective computing power could be substantial. For example, a laptop could handle intensive computational tasks, smartphones could manage coordination and lightweight processing, and desktop computers could contribute their onboard computing power. Meanwhile, other IoT devices such as smart speakers, security cameras, and vehicles could serve as data sources, providing valuable real-world inputs like voice commands, visual feeds, and environmental parameters. The language model would effectively run and train across this entire device cluster, leveraging both computing power and diverse training data from the environment. This distributed execution requires new frameworks that can intelligently decompose and distribute model computations based on each device’s capabilities and current load. The system must dynamically balance workloads - when the security cameras are idle at night, they could take on additional compute tasks, while during peak usage hours, the load could shift to other devices. This requires innovations in real-time resource allocation, task scheduling across heterogeneous hardware, and efficient inter-device communication protocols to ensure the collective computing power is optimally utilized [98].